OpenStudy (18jonea):

Solve |5x – 3| ≤ 3. @mehek14

OpenStudy (18jonea):

@texaschic101

OpenStudy (texaschic101):

| 5x - 3 | <= 3 5x - 3 <= 3 and 5x - 3 >= -3 these are the 2 inequalities that you will have to solve to arrive at your answers

OpenStudy (texaschic101):

read and learn... http://www.purplemath.com/modules/absineq.htm

OpenStudy (18jonea):

so would it be 0<= x<= 6/5 @texaschic101

OpenStudy (anonymous):

Yes, your answer is right @18jonea